Solve ∛(5x+8)+3=4


Answer by unlockmath(1688)

Hello,
Let's do this:
∛(5x+8)+3=4
Subtract 3 from both sides to get:
∛(5x+8)=1
Do each side to the 3rd power:
[∛(5x+8)]^3=1^3
Will give us:
5x+8=1
Now subtract 8 from both sides:
5x=-7
and divide by 5 to get:
x=-7/5
There we go.
Make sense?
RJ